OpenStudy (anonymous):

3(x+1=-2(x-1)+6 Solve for x

OpenStudy (anonymous):

3x + 3 = -2x +2 + 6 -3 - 3 3x = -2x + 2 +3 3x = -2x + 5 +2x + 2x 3x + 2x = 5 5x = 5 /5 /5 x = 1

OpenStudy (anonymous):

3x+3=-2 -3 -3 3x=-5 -5*x=-5x -5*1=-5 -5x- -5 +6 +5 +5 -5x=11 -5x divided by -5 11 divided by -5 X=-2.2

OpenStudy (anonymous):

i think it's 1 .... 3(x+1)=-2(x-1)+6 3(1+1) = -2(1-1)+6 3 + 3 = -2 + 2 + 6 6 = 6